Joe has a collection of nickels and dimes that is worth $2.15. If the number of dimes was doubled and the number of nickels was increased by 28, the value of the coins would be $4.65. How many nickels and dimes does he have?
Answer by mananth(16949)

Nickels -------x
dimes ----------y
5x+10y =215...................(1)
Nickels------x+28
Dimes---------2y
5(x+28)+10*2y= 465
5x+140+20y=465
5x+20y=325--------------------(2)

subtract (2) from (1)
-10y=-110
/-10
y=11 dimes
plug y in (1) to get value of x
x=21 nickels
CHECK
21*5+10*11=215